OpenStudy (anonymous):

to find limit of this polynomial x->0, ( sqrt(1+tan x)-sqrt(1+sin x) )/x^3, can i just say, as x ->0, sqrt(1+tan x) = 1, and sqrt(1+sin x)=1 so the numerator will become 0?

OpenStudy (psymon):

Yeah, but it looks like you have an x^3 in the denominator, which would just make it undefined as it goes to 0.

OpenStudy (psymon):

\[\frac{ \sqrt{1+tanx}-\sqrt{1+sinx} }{ x ^{3} }\] Is this correct?

OpenStudy (anonymous):

not clear, but it sure as hell is not a polynomial

OpenStudy (anonymous):

i think the limit is \(\frac{1}{4}\) power series might do it, or maybe l'hopital, but that will be a drag

OpenStudy (anonymous):

Yeah psymon